OpenStudy (anonymous):

what is the indefinite integral of 1/x^3

hartnn (hartnn):

1/x^3= x^-3.....integral x^n is (x^(n+1))/(n+1)

OpenStudy (turingtest):

just to make pretty what hartnn wrote\[\large\frac1{x^3}=x^{-3}\]and\[\large\int x^ndx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

thanks ...i've got it

OpenStudy (turingtest):

welcome

OpenStudy (anonymous):

what about the indefinite integral of y^3(2y^2-3)

OpenStudy (turingtest):

distribute, then apply the same rule but please post each Q separately thanks

hartnn (hartnn):

multiply y^3 inside.....which gives 2y^5-3y^3...then apply same formula to both terms